OpenStudy (anonymous):

Solve for y, then find the value of y when given x = –2. 6x = 7 – 4y A. -12 B. 7/4 C. 19/4 D. 24

OpenStudy (anonymous):

just place -2 in x and solve

OpenStudy (anonymous):

6(-2) = 7 - 4y now solve for y

OpenStudy (shaik0124):

4y=7-6x 4y=7-6(-2) 4y=7+12 4y=19 y=19/4

OpenStudy (anonymous):

Thank for the help guys!